OpenStudy (rock_mit182):

A man purchases a $2000 diamond ring for a down payment of $200 and monthly installments of $88 for 2 years. Assuming that interest is compounded monthly, what interest rate is he paying?

OpenStudy (anonymous):

.0627 or 6.27% annually

OpenStudy (rock_mit182):

well is awesome "the answer" but i need more step through the problem

OpenStudy (rock_mit182):

could anybody explain to me how to solve this ?

OpenStudy (radar):

Do you know the compounding interest formula? That might be a good starting point. The formula I have is A = P(1 + r/n)^(nt) where n would be 12 for this problem. You would need to solve for r

OpenStudy (rock_mit182):

n is always the amount of time ?

OpenStudy (radar):

n is the number of compoundings per year t is the number of years in this case t = 2

OpenStudy (rock_mit182):

p = principle amount ?

OpenStudy (radar):

n=12 t=2 P= principal A= Balance

OpenStudy (radar):

You may want to use a google and find a financial calculator. What kind of formula does your reference text suggest?

OpenStudy (rock_mit182):

\[A = 2*10^{3}(1+ \frac{ r }{ 12 })^{12*2} \]

OpenStudy (rock_mit182):

well there are three

OpenStudy (rock_mit182):

If a loan Ap is to be repaid in n regular equal payments with interest rate i per time period, then the size R of each payment is given by

OpenStudy (rock_mit182):

\[R =\frac{ iA _{p} }{ 1-(1+i)^{-n} }\]

OpenStudy (rock_mit182):

i can use that one ?

OpenStudy (anonymous):

that's the formula i used, however that's annual payment so u have to divide the r by 12, solving for i is always tricky algebraically so i just use excel

OpenStudy (radar):

Don't forget he did make a 200 dollar down payment and the amount would be $1800.00

OpenStudy (rock_mit182):

well let me see

OpenStudy (rock_mit182):

\[\frac{ 1- (1+i)^{-n}R }{ A } =i\]

OpenStudy (rock_mit182):

\[\frac{ 1-R }{a(1+i)^{n} } = i \]

OpenStudy (radar):

The calculator I found states the interest rate is 15.845% Need to look further into that.

OpenStudy (rock_mit182):

i guess i did something wrong already

OpenStudy (anonymous):

radar did u use pmt function on excel?

OpenStudy (radar):

Well it might be that when you consider 24 payments of $88 is $2,112.00 which is $312.00 of interest over a 24 month period.

OpenStudy (anonymous):

right but the payments are discounted at different times since you pay monthly

OpenStudy (radar):

No I am using this calculator, I will send the link http://www.mycalculators.com/ca/loancalcm.html

OpenStudy (rock_mit182):

Can I calculate this on wolfram alpha ?

OpenStudy (anonymous):

yeesh im a moron i thought the payment was 80 not 88

OpenStudy (anonymous):

radar's answer is right

OpenStudy (radar):

The original formula I provided a balance but not the interest rate when the monthly payment is given. So it needs to be modified somehow that when the payment is given you can figure out the interest rate. I need to ponder this....

OpenStudy (rock_mit182):

guys the true I'm not good at all on this kind of math

OpenStudy (rock_mit182):

I thought that i could figure out this by myself not with a calculator

OpenStudy (rock_mit182):

I thought that i could figuere out this by myself not with a calculator

OpenStudy (rock_mit182):

Anyway thanks a lot! to both of you